OpenStudy (anonymous):

The sum of two consecutive integers is at least 36. What is the least possible pair of integers? (Points : 1) 16 and 17 17 and 18 18 and 19 19 and 20

OpenStudy (anonymous):

"At least' means : It can be that number and greater than that number too, \( \ge\)..

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

Let the first number be \(x\), then its consecutive will be : \(x+1\)

OpenStudy (anonymous):

So, according to question, their sum should be at least 36 ie : \(x + (x+1) \ge 36\) \(2x + 1 \ge 36\) \(2x \ge 36\) \(x \ge 18\)..

OpenStudy (anonymous):

so b?

OpenStudy (anonymous):

@waterineyes

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

yes @Nnesha here :)))

Nnesha (nnesha):

waterineyes already solved that and like he said this sign mean greater than and equal to so x >= 18 and other number should be equal or it should be greater than 18 so which one should be

OpenStudy (anonymous):

b?

Nnesha (nnesha):

17 is bigger than 18 ???

OpenStudy (anonymous):

no, c

Nnesha (nnesha):

yep right :)

OpenStudy (anonymous):

thank you so so much

Nnesha (nnesha):

my pleasure :) :)

Nnesha (nnesha):

thanks to @waterineyes too thanks water

OpenStudy (anonymous):

Sorry, Electricity was not there and I was operating my pc on UPS.. And after some time UPS said, "I have given enough time to you for shut down, but now I will do it my way" :P

OpenStudy (anonymous):

Question is : What is the least possible pair of integers? You should be careful that "least possible pair" you have to find, it means the minimum possible value of \(x\), and as our \(x \ge 18\), In words, x can be \(18\) or greater than \(18\).. So, that is why we are taking \(x\) as \(18\) which is its least value.. Hence, \(18\) and \(19\)..

OpenStudy (anonymous):

Strictly speaking, if in question, you find no word as "least possible pair", then \(c\) and \(d\), both will be your answer.. because \(d\) also satisfies the condition.. So, remember the difference.. :)

OpenStudy (anonymous):

Thank @Nnesha for taking over... :)

OpenStudy (anonymous):

*Thanks.. :)

Nnesha (nnesha):

no problem its my pleasure :) @waterineyes